# Comprehensive Evaluation on Lexical Normalization: Boundary-Aware Approaches for Unsegmented Languages

### v1

@@ 7.4 Investigation of Training Data Size
Second, in model-specific comparisons, precision and recall follow different patterns. Precision shows no clear differences across models. In contrast, recall is consistently highest for Sarashina2-7b, followed by Sarashina2.2-1b, and lower for DeBERTa and T5; this indicates that the Sarashina-series models generalize well in terms of coverage, even when the training data size is small.
Figure 2: JMLN test results for each training data size.
@@ 7.5 Results Across Domains
As shown in Table 6, we evaluated performance of selected models—DeBERTa-large, T5-large (Span), and Sarashina-2.2-3B (Span)—for each domain test set of JMLN.
First, to explore what makes a domain difficult, we examined an indicator: the proportion of non-standard surface tokens in the test set that are not found among the training set’s non-standard surfaces (Surf-Outside-Train rate). We then computed the Pearson correlation coefficients $r$ between the indicator and the average F0.5 scores across the three models, obtaining a strong negative correlation ($r=-0.78$). Notably, for 9 out of 10 domains with a Surf-Outside-Train rate below 0.5 had average F0.5 scores above 0.7, whereas all 4 domains with a rate above 0.5 had average scores below 0.7.
@@
Table 7: Models’ prediction statistics on JLMN dev set (DeBERTa: large, T5: large, Sarashina: 3B).
7.6 Error Analysis
For the three models evaluated in §7.5, we analyzed output patterns on the JMLN dev set. Specifically, we counted predictions, true positives (TP), false positives (FP), and false negatives (FN) for each model, and measured (i) the proportion of predicted normalized forms appearing in the UniDic lexicon (Norm-In-Lex rate) and (ii) the proportion of original surface forms for TP, FP, and FN instances that matched any non-standard forms in the training set (Surf-In-Train rate). Table 7 shows the results averaged over two runs per model.
DeBERTa exhibited a notably lower Norm-In-Lex rate than the other models (0.921–0.928) and the gold standard (0.947). By manually inspecting error cases, we found that DeBERTa’s restored tokens within spans sometimes formed nonsensical words. [^12] E.g., {CJK}UTF8ipxmかど (orig. {CJK}UTF8ipxmヶド, gold {CJK}UTF8ipxmけど/{CJK}UTF8ipxmけれど, “but”) and {CJK}UTF8ipxmだじりん (orig. {CJK}UTF8ipxmだぁりん, gold {CJK}UTF8ipxmダーリン, “darling”). Both suggest that the model’s independent prediction at each MASK position makes it especially prone to such errors.
Regarding the Surf-In-Train rate, all models exhibited similar trends. For TPs, approximately 70–80% of the original surface forms were known (i.e., appeared in the training set), indicating that many correct predictions relied on the seen normalization instances. For both FPs and FNs, only approximately 20–30% were known, indicating that the majority of errors involved unseen expressions. This suggests considerable room for improving generalization in normalizing unseen cases. These results align with the findings in §7.5.
@@ 8 Conclusion
An state-of-the-art encoder-only model, ModernBERT Warner et al. (2024), might achieve performance on par with or exceeding the models we evaluated. Due to computational and time constraints, this remains future investigation.
@@
Our word-level normalization metrics treat any span mismatch as an error—even if the predicted normalization is semantically valid—which we observed especially in decoder-only model outputs (see Appendix D.5 for examples). Such span differences have little impact on most downstream applications, so the practical usefulness of the outputs may exceed the scores reported. To complement word-level metrics, we also provide sentence-level exact-match accuracy and the chrF score Popović (2015) in Appendix D.2 and D.2.
@@
MeCab is available under GPL, LGPL, and BSD License. UniDic (“unidic-cwj-3.1.0”) is available under GPL v2.0, LGPL v2.0, and New BSD License. sacreBLEU is available under Apache License 2.0 (we used this software for preliminary experiments shown in Appendix D). The licenses for the datasets and pre-trained models are listed in Appendices C.2 and A.2 (Table 19). Our use of these resources for academic research aligns with their intended use. We will release our JMLN dataset for academic research in information science; it will include only annotation information and not the original texts.
@@ Appendix A JMLN Dataset
A.1 Data Statistics
The detailed dataset statistics are shown in Table 8 and Table 9.
@@ 14 SK 684 11,935 286
Table 8: Statistics of the JMLN Cur-sets.
@@ 14 SK 200 2,182 16
Table 9: Statistics of the JMLN Rand-sets.
A.2 Data Sources and Licenses
@@ 12 JW Japanese Wikipedia Typo Dataset Encyclopedia edit history –2021
A.3 Word Category Definition
@@
A.4 Standard/Non-Sandard From Definition
@@
Example Standard forms Translation
Vocabulary type Neologisms/Slang {CJK}UTF8ipxmコピペ copy and paste
Proper names {CJK}UTF8ipxmドラクエ Dragon Quest
Onomatopoeia {CJK}UTF8ipxmキラキラ glitter
Interjections {CJK}UTF8ipxmおお oops
Dialect words {CJK}UTF8ipxmほんま truly
Foreign words {CJK}UTF8ipxmEASY easy
Ancient words* {CJK}UTF8ipxm[行く] べし should [go]
[…]
Emoticons/AA {CJK}UTF8ipxm(^−^)
Variant-form type Character type variants {CJK}UTF8ipxmカワイイ {CJK}UTF8ipxmかわいい,可愛い cute
Alternative representations {CJK}UTF8ipxm大きぃ {CJK}UTF8ipxm大きい big
Sound change variants {CJK}UTF8ipxmおいしーい {CJK}UTF8ipxmおいしい,美味しい tasty
Typographical errors {CJK}UTF8ipxmつたい {CJK}UTF8ipxmつらい,辛い tough
Table 11: Word categories extended from Higashiyama et al. (2021b). New categories are marked with “*.” “[]” indicates the context. {CJK}UTF8ipxm[行く] にゃ” is a kitten-style sentence ending and might be expressed as “[I go,] meow” or “[Goin’]nya.” {CJK}UTF8ipxmおはこんばんにちは” is a coined blend of “good morning,” “good afternoon,” and “good evening,” and might be expressed as “Good morn-noon-evening.”
@@
Sent Word Cate VForm Word POS Lem Cate Cateb VForm VFormb
@@
Table 12: Statistics on inter-annotator agreement. The “Number” columns show counts of annotated sentences (Sent), words, word categories (Cate), variant forms (VForm) for each annotator (formatted as “value1/value2”). The “F1 agreement score” columns report F1 scores for word segmentation (Seg), parts-of-speech (POS), lemmas (Lem), Cate, and VForm. Binary agreement on the presence or absence of a category/variant-form assignment is indicated by “$b$.”
A.5 Standard Form Annotation
The annotation process first identifies word boundaries and then assigns each word to (a category and) a standard form ID, as in the example below (meaning “Super tired today.”); for simplicity, the example displays the standard-form string instead of the ID. This approach makes it possible to obtain explicit mappings between non-standard words and their standard forms.
Original text {CJK}UTF8ipxm今日はスゴクツカレタ~
Segmented text {CJK}UTF8ipxm今日 |||| は |||| スゴク |||| ツカレタ~
Standard forms {CJK}UTF8ipxm   –    ||||  –  |||| すごく |||| 疲れた
However, when only sentence-level normalization is provided, aligning standard and non-standard forms is not straightforward. In our example, the contiguous span mapping from “{CJK}UTF8ipxmスゴクツカレタ~” to “{CJK}UTF8ipxmすごく疲れた” (“super tired”) is easily identified, but determining precise word-level alignments remains challenging.
A.6 Inter-Annotator Agreement
@@
Appendix B Encoder-based Approach Variants
In addition to the Full-Seg approach, we introduce two variant approaches based on the encoder-base detect&infill method: Part-Seg and Full-Seg-POS. Experimental results on these variants are reported in D.2.
@@ Appendix B Encoder-based Approach Variants
Appendix C Detailed Experimental Settings
C.1 Definition of Evaluation Metrics
@@ C.3 Model Hyperparameters
Warmup ratio 0.1
@@
Warmup ratio 0.1
@@
C.4 Prompts for Decoder-only Models
@@
C.5 Computational Budget for Fine-tuning
@@
C.6 Throughput Calculation Setting
@@
C.7 Word Coverage Indicators
@@
Appendix D Detailed Experimental Results
D.1 Accuracy Across Model Series
@@
D.2 Comparison of Encoder Approaches
As a preliminary experiment, we compared variants of encoder-based methods—Part-Seg, Full-Seg, and Full-Seg-POS, introduced in Appendix B—on the JMLN dev set. We also report the F1 scores for word segmentation, POS tagging, and length prediction (Lenp: tokens in positive non-standard words; Lenn: other tokens), as well as additional normalization metrics: sentence-level exact match accuracy (S-Accp: accuracy for sentences containing at least one positive non-standard word; S-Accn: accuracy for other sentences) and the chrF score Popović (2015) implemented in sacreBLEU Post (2018). [^21] We calculated chrF scores with the default options (signature: nrefs:1|case:mixed|eff:yes|nc:6|nw:0| space:no|version:2.5.1). Table 16 shows the results.
The encoder-only models (3 backbone models×$\times$2 model size) with the Full-Seg approach obtained +0.01–0.05 F0.5 gains from the Part-Seg counterpart, indicating the importance of learning full word segmentation tasks. Adding POS tagging task showed no clear improvements.
D.3 Comparison of Generative Approaches

### v2

@@ 7.4 Investigation of Training Data Size
Second, in model‐specific comparisons, precision and recall follow different patterns. Precision shows no clear differences across models. In contrast, recall is consistently highest for Sarashina2-7B, followed by Sarashina2.2-1B, and lower for DeBERTa and T5; this indicates that the Sarashina-series models generalize well in terms of coverage, even when the training data size is small.
@@ 7.5 Results Across Domains
As shown in Table 6, we evaluated the performance of selected models—DeBERTa-large, T5-large (Span), and Sarashina-2.2-3B (Span)—on each domain test set of JMLN (results on grouped domains are provided in Appendix E.2).
First, to explore what makes a domain difficult, we examined an indicator: the proportion of non-standard surface tokens in the test set that are not found among the training set’s non-standard surfaces (Surf-Outside-Train rate). We then computed the Pearson correlation coefficients $r$ between the indicator and the average F0.5 scores across the three models, obtaining a strong negative correlation (r=0.78r=-0.78). Notably, for 9 out of 10 domains with a Surf-Outside-Train rate below 0.5 had average F0.5 scores above 0.7, whereas all 4 domains with a rate above 0.5 had average scores below 0.7.
@@
7.6 Informal Span Detection Accuracy
We evaluated the performance of the informal word span detection task for the same three Japanese models as in §7.5.131313We report the detection accuracy of the Thai models in Appendix E.5. In this task, a model’s prediction is regarded as correct if the predicted span of a non-standard word matches the gold standard, irrespective of the predicted standard form. Table 7 shows the detection results, the normalization results, and the precision and recall differences between the two tasks.
DeBERTa achieved the best precision and F0.5 score among the models for the detection task, but its precision and recall dropped by about 0.09–0.12 for the normalization task. Sarashina2.2 achieved the best recall for the detection task, with smaller drops of about 0.04–0.05 precision and recall for the normalization task, resulting in the best overall normalization performance. T5 yielded the worst results; the precision/recall differences between the two tasks were intermediate between those of the remaining models. These results suggest that DeBERTa and T5 leave notable room for improvement in generating accurate standard forms, whereas the best-performing Sarashina2.2 shows only a minor gap. Thus, further accuracy gains will likely depend on better handling of cases where span detection fails.
7.7 Error Analysis
For the three models evaluated in §7.5 and §7.6, we analyzed output patterns on the JMLN dev set. Specifically, we counted predictions, true positives (TP), false positives (FP), and false negatives (FN) for each model, and measured (i) the proportion of predicted normalized forms appearing in the UniDic lexicon (Norm-In-Lex rate) and (ii) the proportion of original surface forms for TP, FP, and FN instances that matched any non-standard forms in the training set (Surf-In-Train rate). Table 8 shows the results averaged over two runs per model.
DeBERTa exhibited a notably lower Norm-In-Lex rate than the other models (0.921–0.928) and the gold standard (0.947). By manually inspecting error cases, we found that DeBERTa’s restored tokens within spans sometimes formed nonsensical words. [^14] E.g., pred. かど, orig. ヶド, gold けど/けれど (“but”) and pred. だじりん, orig. だぁりん, gold ダーリン (“darling”). Both suggest that the model’s independent prediction at each MASK position makes it especially prone to such errors.
Regarding the Surf-In-Train rate, all models exhibited similar trends. For TPs, approximately 70–80% of the original surface forms were known (i.e., appeared in the training set), indicating that many correct predictions relied on the seen normalization instances. For both FPs and FNs, only approximately 20–30% were known, indicating that the majority of errors involved unseen expressions. This suggests considerable room for improving generalization in normalizing unseen cases. These results align with the findings in §7.5. For actual prediction examples, see Appendix E.7.
@@ 7.7 Error Analysis
Table 8: Models’ prediction statistics on JLMN dev set (DeBERTa: large, T5: large, Sarashina2.2: 3B).
@@ 8 Conclusion
Generalizability of Experimental Results
Our findings and conclusions are presented in the context of the datasets, languages, and models used in this study.
@@
A state-of-the-art encoder-only model, ModernBERT Warner et al. (2024), might achieve performance on par with or exceeding the models we evaluated. Due to computational and time constraints, this remains future investigation.
@@
Our word-level normalization metrics treat any span mismatch as an error—even if the predicted normalization is semantically valid—which we observed especially in decoder-only model outputs (see Appendix E.7 for examples). Such span differences have little impact on most downstream applications, so the practical usefulness of the outputs may exceed the scores reported. To complement word-level metrics, we also provide sentence-level exact-match accuracy and the chrF score Popović (2015) in Appendix E.3 and E.4.
@@
MeCab is available under GPL, LGPL, and BSD License. UniDic (unidic-cwj-3.1.0) is available under GPL v2.0, LGPL v2.0, and New BSD License. sacreBLEU is available under Apache License 2.0 (we used this software for preliminary experiments shown in Appendix E). The licenses for the datasets and pre-trained models are listed in Appendices D.2 and A.3 (Table 23). Our use of these resources for academic research aligns with their intended use. We will release our JMLN dataset for academic research in information science; it will include only annotation information and not the original texts. [^15] The original texts have already been released by the original dataset providers, and we will publish our annotation dataset along with a tool that integrates those texts with the annotations, enabling users to reconstruct the complete dataset.
@@
Acknowledgments
We would like to thank the anonymous reviewers and meta reviewers at ACL Rolling Review for their constructive feedback, especially the reviewer whose detailed comments and extensive suggestions were helpful in improving the quality of this paper. To construct our dataset, we made use of existing datasets: the Balanced Corpus of Contemporary Written Japanese, the Rakuten Dataset, the Multilingual Amazon Reviews Corpus, the Japanese Wikipedia Typo Dataset, the Nagoya University Conversation Corpus, and the Japanese and Chinese Skype Conversation Corpus, as well as the Recruit Dataset provided by Recruit Co., Ltd. and the Niconico Dataset provided by DWANGO Co., Ltd., both through the IDR Dataset Service of the National Institute of Informatics.
@@ Appendix A JMLN Dataset
A.1 Details on Data Creation
Before conducting the sentence selection step described in §3, we divided each candidate set from each of the original 14 datasets into two subsets: a manually curated (Cur) set and a random (Rand) set. The successive steps for the Cur set are those described in §3, and the resulting annotation data are divided into train, dev, and testc sets, as in §6.1.
For the Rand set, annotators sequentially selected sentences from candidates of each original dataset, provided that they did not contain ethically problematic contents or unclear meaning. We consider the entire resulting annotation data as an additional test set, which we refer to as the testr set. This dataset is considered to exhibit a distribution of non-standard word frequencies close to that of natural text, making it useful for evaluations that assume such data. However, experiments using this dataset are left for future work.
A.2 Data Statistics
The detailed dataset statistics of JMLN are shown in Table 9 and Table 10. Note that sentences in the Cur set are divided into train, dev, and testc sets, and the Rand set corresponds to the testr set.
@@ 14 SK 684 11,935 286
Table 9: Statistics of the JMLN Cur set.
@@ 14 SK 200 2,182 16
Table 10: Statistics of the JMLN Rand set.
A.3 Data Sources and Licenses
@@ 12 JW Japanese Wikipedia Typo Dataset Encyclopedia edit history –2021
Example Standard forms Translation
Vocabulary type Neologisms/Slang コピペ copy and paste
Proper names ドラクエ Dragon Quest
Onomatopoeia キラキラ glitter
Interjections $お$ oops
Dialect words ほんま truly
Foreign words EASY easy
Ancient words* [行く] べし should [go]
Character endings* [行く] にゃ $\dagger$
Blend words* おはこんばんにちは $\ddagger$
Emoticons/AA (^−^)
Variant-form type Character type variants カワイイ かわいい,可愛い cute
Alternative representations 大きぃ 大きい big
Sound change variants おいしーい おいしい,美味しい tasty
Typographical errors つたい つらい,辛い tough
Table 12: Word categories extended from Higashiyama et al. (2021b). New categories are marked with “*.” “[]” indicates the context. “[行く] にゃ” is a kitten-style sentence ending and might be expressed as “[I go,] meow” or “[Goin’]nya.” “おはこんばんにちは” is a coined blend of “good morning,” “good afternoon,” and “good evening,” and might be expressed as “Good morn-noon-evening.”
A.4 Word Category Definition
@@
A.5 Standard/Non-Sandard From Definition
@@
Sent Word Cate VFCate Word POS Lem Cate Cateb VFCate VFCateb
@@
Table 13: Statistics on inter-annotator agreement. The “Number” columns show counts of annotated sentences (Sent), words, word categories (Cate), variant-form type categories (VFCate) for each annotator (formatted as “value1/value2”). The “F1 agreement score” columns report F1 scores for word segmentation (Seg), parts-of-speech (POS), lemmas (Lem), Cate, and VFCate. Binary agreement on the presence or absence of a category assignment is indicated by “bb.”
A.6 Standard Form Annotation
The annotation process first identifies word boundaries and then assigns each word to (a category and) a standard form ID, as in the example in Table 14; for simplicity, the example displays the standard-form string instead of the ID. This approach makes it possible to obtain explicit mappings between non-standard words and their standard forms.
However, when only sentence-level normalization is provided, aligning standard and non-standard forms is not straightforward. In our example, the contiguous span mapping from “スッゴクツカレタ~” to “すごく疲れた” (“super tired”) is easily identified, but determining precise word-level alignments remains challenging.
Original text 今日はスッゴクツカレタ~
Segmented text 今日 $|$ は $|$ スッゴク $|$ ツカレタ~
Standard forms    –    $|$  –  $|$ すごく $|$ 疲れた
Table 14: An example sentence and its annotation information. The original text means “Super tired today.”
A.7 Inter-Annotator Agreement
@@
Appendix B Discussion on the Task Definition and Alternative Approaches
As discussed in §4, we adopted a boundary-aware LN task with explicit span identification, which enables direct span-level evaluation. A simpler text-to-text conversion task can instead be evaluated with sentence-level metrics. However, such metrics (e.g., chrF) allow only relative comparisons: sentences with few normalization targets may achieve high scores even for a trivial “leave-as-is” baseline, indicating that these scores do not reflect absolute output quality.
Moreover, tokenizers used in Japanese encoder-only models may alter characters—for example, splitting a single-character ellipsis “…” (U+2026) into three consecutive period characters “.” (U+002E)—causing sentence-level scores to fluctuate even when those characters lie outside normalization spans. Hence, selecting a fair ground-truth sentence—whether the original or a tokenizer-converted sentence—across models is non-trivial. These considerations impose constraints on character-by-character transduction systems.
One might then ask whether the use of an off-the-shelf word segmenter removes the need to introduce a word boundary prediction step in the encoder-based infilling approach, while still maintaining the feasibility of span-level evaluation. Possible alternatives using an external word segmenter include pre-normalization segmentation and post-normalization segmentation.
On the one hand, the former approach segments the original input text with a segmenter and then normalizes, for example, at the WordPiece level. This preserves word-boundary positions through prefix symbols (e.g., “ye” and “##a” for “yea”), similar to Muller et al.’s method for English LN. However, off-the-shelf word segmenters trained on canonical text often produce segmentation errors on informal words, which can directly propagate to the normalization results.
On the other hand, the latter approach segments the normalizer’s outputs with a segmenter, which generally avoids salient errors when the outputs are accurate. Nevertheless, span-level evaluation under this setting requires non-trivial alignment between ad-hoc word boundaries in the output and those in the original text (see an example in Appendix A.6).
Appendix C Encoder-based Approach Variants
In addition to the Full-Seg approach, we introduce two variant approaches based on the encoder-base detect-and-infill method: Part-Seg and Full-Seg-POS. Experimental results on these variants are reported in E.3.
@@ Appendix C Encoder-based Approach Variants
Appendix D Detailed Experimental Settings
D.1 Definition of Evaluation Metrics
@@ D.2 Pretrained Models
Warmup steps 1 training epoch
@@
Warmup steps 1 training epoch
@@
D.3 Model Hyperparameters
D.4 Prompts for Decoder-only Models
@@
D.5 Computational Budget for Fine-tuning
@@
D.6 Throughput Calculation Setting
@@
D.7 Word Coverage Indicators
@@
Data IDs Register DeBERTa-L T5-L Sarashina2.2-3B
P R P R P R
01 Q&A site 0.654 0.416 0.572 0.369 0.717 0.578
02–03 Blog 0.750 0.595 0.761 0.591 0.802 0.685
04–06,08 Reviews 0.837 0.630 0.769 0.568 0.808 0.687
07 Recipes 0.815 0.684 0.817 0.663 0.834 0.734
09 Video desc. 0.627 0.538 0.666 0.492 0.730 0.629
10 Forum 0.850 0.747 0.783 0.677 0.808 0.772
11 Social media 0.613 0.511 0.600 0.464 0.763 0.677
12 Wiki hist. 0.528 0.140 0.197 0.095 0.509 0.280
13–14 Conv. trans. 0.816 0.662 0.699 0.537 0.812 0.710
Table 17: JMLN test results of representative three models for each aggregated domain.
@@
Appendix E Detailed Experimental Results
E.1 Accuracy Across Model Series
@@
E.2 Results Across Grouped Domains
Table 17 shows the performance of three selected models as in §7.5—DeBERTa-large, T5-large (Span), and Sarashina-2.2-3B (Span)—on the grouped domain test set of JMLN, in which domains of a similar type from different data sources (e.g., all blog sites) are combined into a single domain group. The results are consistent with those on the single-domain test sets. Sarashina-2.2-3B achieved the highest recall across all grouped domains and also obtains the highest precision in 5 out of 9 groups, while DeBERTa-large achieved the highest precision in the remaining groups.
E.3 Comparison of Encoder Approaches
As a preliminary experiment, we compared variants of encoder-based methods—Part-Seg, Full-Seg, and Full-Seg-POS, introduced in Appendix C—on the JMLN dev set. We also report the F1 scores for word segmentation, POS tagging, and length prediction (Lenp: tokens in positive non-standard words; Lenn: other tokens), as well as additional normalization metrics: sentence-level exact match accuracy (S-Accp: accuracy for sentences containing at least one positive non-standard word; S-Accn: accuracy for other sentences) and the chrF score Popović (2015) implemented in sacreBLEU Post (2018). [^24] We calculated chrF scores with the default options (signature: nrefs:1|case:mixed|eff:yes|nc:6|nw:0| space:no|version:2.5.1). Table 20 shows the results.
[…]
Backbone Approach Det Norm Det-Norm
P R F0.5 P R F0.5 P R
RoBERTa-base Full-Seg 0.804 0.596 0.751 0.713 0.529 0.666 0.091 0.067
T5-base Span 0.715 0.684 0.708 0.645 0.618 0.640 0.069 0.066
mT5-large Span 0.725 0.670 0.713 0.660 0.609 0.649 0.066 0.061
Llama3.2-Typhoon2-3b Span 0.717 0.731 0.720 0.656 0.668 0.658 0.062 0.063
Llama3.2-Typhoon2-8b Span 0.722 0.741 0.726 0.661 0.678 0.664 0.061 0.063
Table 18: VISTEC test results of selected Thai models for detection (det) and normalization (norm) tasks. The best score among compared systems is highlighted in bold.
E.4 Comparison of Generative Approaches